Order: Itraconazole (Sporanox) 4 mg/kg/d divided in equal dosages q.8h. The multi-dose vial label reads 10 mg/ml. How many
Answer #1

The required dose = 4mg/ kg/day=340kg

The total dose = 10mg /ml

So,the dose from total dose = 340/10*1=34ml to be given for 24hours

So, q8h =34/3= 11.3= 11ml

So, the answer is 10ml (nearest tenth)